Suppose that Fortunate Ventures, Inc.'s, common shares are currently selling for $23 a share with 7 million shares outstanding. The firm also has 12,000 bonds outstanding, which are selling at 96 percent of par. If Fortunate Ventures was considering an active change to its capital structure to have a D/E ratio of 0.5, which type of security (stock or bonds) would the firm need to sell to accomplish this, and how much would it have to sell?

A. Bonds; $54,325,000

B. Stock; $38,460,000

C. Bonds; $68,980,000

D. Stock; $44,698,000
Answer #1

Debt/Equity=(12000*96%*1000+y)/(23*7*10^6)

=>(12000*96%*1000+y)/(23*7*10^6)=0.5
=>-12000*96%*1000+0.5*23*7*10^6=y
=>y=68980000

So, Bonds=68980000